Arieonna:

Round 5969.03 to the nearest hundreth

Arieonna:

@lars can u help me

Arieonna:

this is the screen shot the number below i written is what i got i need to figure out what the nearest hundredth is

K1NGofPadlet:

@Arieonna Hey, never fear, your King is here: Rounding 5969.03 to the most proximate hundredth would be 5969.03 rounded to two decimal places, which is 5969.03.

Arieonna:

@k1ngofpadlet wrote:
@Arieonna Hey, never fear, your King is here: Rounding 5969.03 to the most proximate hundredth would be 5969.03 rounded to two decimal places, which is 5969.03.
but i thought it would be 5969 instead of the answer i got doing my math

Arieonna:

?

K1NGofPadlet:

Seems to me that there's no further digit to round it to, Arieonna.

Arieonna:

o ok so if their mite not do u think i did the math wrong?

K1NGofPadlet:

@arieonna wrote:
o ok so if their mite not do u think i did the math wrong?
Let me redo the whole thing. One sec. Just to see.

K1NGofPadlet:

The formula for the volume of an oblique cylinder is V = Ah, where A is the area of the base and h is the height. To find the area of the base, we require to first find the radius. We can do this by utilizing the Pythagorean theorem: r^2 = a^2 + h^2 r^2 = 10^2 + 19^2 r^2 = 461 r = √461 Now we can find the area of the base: A = πr^2 A = 3.14(√461)^2 A = 3.14(461) A = 1449.94 Conclusively, we can plug in the values for A and h to find the volume: V = Ah V = 1449.94(19) V = 27529.86 Rounding to the most proximate hundredth, the volume of the oblique cylinder with a = 10 cm and h = 19 cm is 27529.86 cubic centimeters.

K1NGofPadlet:

If that doesn't work, then to find the volume of the oblique cylinder with a = 10 cm and h = 19 cm, we can utilize the formula: V = πr^2h First, we require to find the radius (r) of the cylinder. We can utilize the Pythagorean theorem to do this: r = √(a^2 - h^2) = √(10^2 - 19^2) = √(-171) ≈ 0 (since the square root of a negative number is not a genuine number) Since we cannot have a negative radius, it appears that the values of a and h are not possible for an oblique cylinder. Ergo, the volume cannot be tenacious with the given quantifications..

Arieonna:

thats a lot for me to read so i would have to round 27529.86???????

Arieonna:

@k1ngofpadlet

Arieonna:

my answer was 5966

K1NGofPadlet:

@arieonna wrote:
My answer was 5966.
Yeah, I didn't think the entire thing was complete without adding the negative radius with the other variables. I would like to see how your teacher explains that to me.